How do you solve:
3x + y = 8
y= x + 4

Answer by kapilsinghi(68)

3x + y = 8 ------------1
y= x + 4 ------------2
subtituting value of y from equation 2 to 1
3x + x + 4 = 8
4x + 4 = 8
4x = 8 -4
4x = 4
x = 1
Equation 2 says y = x + 4
so y = 1 +4
y = 5
kapilsinghi123@gmail.com